### v2

# MBIAS: Mitigating Bias in Large Language Models While Retaining Context

Abstract
In addressing the critical need for safety in Large Language Models (LLMs), it is crucial to ensure that the outputs are not only safe but also retain their contextual accuracy. Many existing LLMs are safe fine-tuned either with safety demonstrations, or rely only on adversarial testing. While able to get safe outputs, they often risk losing contextual meaning as they mitigate bias and toxicity. In response, we present MBIAS, a LLM framework instruction fine-tuned on a custom dataset specifically designed for safety interventions. MBIAS aims to address the significant issues of bias and toxicity in LLMs generations that typically manifest as underrepresentation or negative portrayals across various demographics, including inappropriate linguistic mentions and biased content in social media. We experiment on MBIAS for safety interventions using various configurations, and demonstrate more than a 30% reduction in overall bias and toxicity while successfully retaining key information. Additionally, a demographic analysis on an out-of-distribution test set confirms the robustness of our approach, with reductions in bias and toxicity exceeding 90% across various demographics. The dataset and instruction fine-tuned MBIAS are made available to the research community  [^1] https://huggingface.co/newsmediabias/MBIAS.
Warning: This paper contains examples that may be offensive or upsetting.
MBIAS: Mitigating Bias in Large Language Models While Retaining Context
Shaina Raza1thanks: Corresponding author: shaina.raza@vectorinstitute.ai, Ananya Raval1, Veronica Chatrath1 shaina.raza, ananya.raval,veronica.chatrath@vectorinstitute.ai 1Vector Institute for Artificial Intelligence
1 Introduction
The rise of generative artificial intelligence (AI) has brought forth substantial ethical and safety challenges, raising concerns of misuse, misinformation, and bias of the generated information Wach et al. (2023). Recent initiatives aim at aligning large language models (LLMs) with ethical norms. These efforts prioritize mitigating biases and enhancing values such as inclusivity, fairness, and non-discrimination Guo et al. (2023). To address harmful, biased, or misleading content (referred to as ‘biased texts’ herein), foundational strategies include implementing safety guardrails in the outputs generated by LLMs. These guardrails provide guidelines and boundaries to ensure AI applications align with ethical standards and societal expectations Att (2024).
Methods such as red-teaming Perez et al. (2022), human feedback during pre-training Casper et al. (2023), and data augmentation methods (e.g., balanced sampling, paraphrasing, or counterfactual data generation) Sadasivan et al. (2023) are often used to reduce biases in LLMs, with the goal of making them safer and more aligned with human values Ganguli et al. (2022); Korbak et al. (2023). In the fine-tuning phase, techniques like instruction tuning, reinforcement learning from human feedback (RLHF), and safety context distillation Ouyang et al. (2022); Qi et al. (2023); Bai et al. (2022) are also used to address unsafe behaviors and improve the model’s ability to generalize.
Despite advancements in the implementation of LLM safety measures, one concern is the loss of actual context or meaning of the original text Schlicht et al. (2024). This delicate balance between two competing goals — reducing biases in the text and preserving informational integrity — highlights a central paradox in bias reduction efforts. Catastrophic forgetting, which occurs when a model forgets previously learned information while acquiring new knowledge, is particularly an issue worth noting while implementing safety mechanisms Luo et al. (2023). Demonstrating the understanding of language models post-safety interventions has thus become a topic of research and discussion Nadeem et al. (2021); Schick et al. (2021).
Recent studies indicate that incorporating even a modest number of safety-focused examples during the instruction-tuning stage can effectively mitigate certain safety concerns Bianchi et al. (2023); Inan et al. (2023); Bai et al. (2022). State-of-the-art LLMs such as GPT-4, Mistral, PaLM, LLaMA-2 and 3, and Claude-2 have been further fine-tuned using high-quality demonstrations, including perturbations and adversarial prompts, to enhance safety. While not enirely foolproof, this safety-tuning enables LLMs to reduce biases in their outputs. Advancing beyond mere demonstrations, we propose instruction fine-tuning a LLM on safety mechanisms. This approach aligns with prior research that emphasizes the importance of the quality and breadth of instruction-tuning data for developing proficient and reliable instruction-following models Touvron et al. (2023); Wang et al. (2024).
Our primary objective in this research is to create a safe LLM that can address bias and toxicity in the outputs while retaining the original knowledge in the message. The primary contributions of this study are as follows:
  • We curate an instruction-tuning dataset with a focus on safety considerations. This dataset comprises paired examples: one containing potentially unsafe elements, such as stereotypes or prejudices, and its corresponding benign (safe) counterpart. The dataset was carefully annotated by human annotators.
  • We propose MBIAS (Mitigating Bias through LLM), an instruction-fine tuned model, built on Mistral2-7B-instruct Jiang et al. (2023). The fine-tuning process involves training the model with our custom dataset that contains examples of both unsafe and safe instances, guiding the model to recognize biases and generate safe responses that retain the meaning of the original input.
  • MBIAS can be adapted for use with other LLMs. For simplicity and ease of implementation, we employ parameter-efficient fine-tuning Ding et al. (2023) to train the model. Similar to Llama Guard Inan et al. (2023), we aim to release model weights with the smallest available model to enhance user-friendliness.
Experimental results on our test set and an out-of-distribution test set across various demographics demonstrate the robustness of our safety interventions in LLMs. We are aware of the ethical implications of modifying user content. However, our aim is to establish a method for fair LLM generations that respects copyright boundaries, while maintaining user trust and autonomy.
2 Related Works
Safety in LLMs Establishing safety measures and protocols is crucial to upholding trust in generative AI and LLMs. Many methods have been proposed to address specific biases in language models. Traditional methods to ensure safe outputs includes examining the embedding space of words and sentences to mitigate biases in the text. Embedding-space-based methods are often applied after training, requiring little-to-no fine-tuning of the model. These methods function as post-processing debiasing steps Liang et al. (2020a); Ungless et al. (2022). Subtraction-based methods are also used to remove biases in language models, which map the embedding space into a neutral one Bolukbasi et al. (2016), maintaining equal distance between non-gendered words and pairs of gendered words. Another method is to ensure that gender-neutral words are orthogonal to the gender direction Zhao et al. (2018). In a related work Zhao et al. (2019), the gendered words are replaced with their opposites in the original training data, and the model is trained on both the original and augmented data. Other methods include efficient fine-tuning for debiasing Gira et al. (2022) and fine-tuning only a small portion of parameters on debiasing data Gira et al. (2022).
Prompt-based debiasing, ranging from prompt-tuning using continuous prompts Yang et al. (2023), to self-supervised zero-shot prompting is also used to detect and reduce bias by controlling model behavior during generation. For example, the Self-Diagnosis and Self-Debiasing methods Schick et al. (2021) reduce the probability of language models generating biased text.
Figure 1: MBIAS architecture showing data preparation and model training with parameter efficient fine tuning.
Debiasing practicality and reliability also depend on dataset selection, evaluation methods, and demographic coverage. Datasets like RedditBias Barikeri et al. (2021), WinoBias Zhao et al. (2019), HolisticBias Smith et al. (2022), RealToxicityPrompts Gehman et al. (2020), and others discussed in Chang et al. (2024) are frequently used for evaluating bias reduction in models. A variety of bias evaluation metrics are used, such as WEAT Liang et al. (2020b), Perspective API API (2024), StereoSet Stereotype Score Nadeem et al. (2021), and LLM alignment methods Wang et al. (2023); Chang et al. (2024).
Despite rapid adoption, safety concerns remain Zhou et al. (2024); Gudibande et al. (2023), particularly around production integration. Most recent LLMs, such as GPT-4, Mistral, PaLM, and Llama-2 and 3, have been instruction fine-tuned with high-quality instruction-following data. While debates persist regarding the competitiveness of finely-tuned instruction models Zhou et al. (2024); Gudibande et al. (2023), there has been rapid adoption of these models among the NLP community.
Numerous considerations on their adoption and integration into production settings are currently under observation. Our focus lies in safety implications for instruction-tuned models: specifically, how these models respond safely to biased instructions, while retaining context. Recent research indicates that training LLMs on demographically perturbed data results in fairer models Qian et al. (2022). We explore whether training on unsafe-benign text pairs achieves better fairness in LLMs, while retaining knowledge.
3 Method
For the purposes of this study, we define bias as any textual content that portrays unfair, inaccurate, or derogatory assumptions about individuals or groups based on their inherent characteristics such as race, gender, ethnicity, or age Raza et al. (2024). The MBIAS architecture can be seen in Figure 1.
3.1 Dataset Preparation
The dataset for this study, extracted from our vast collection of about 3.7M records [^2] Anonymous due to double blind review., spans diverse content from news and social media platforms. This dataset is entirely in English and covers 200+ safety risk aspects. We chose a statistically significant subset of 8.5k records for this work, balancing diversity with computational efficiency, ensuring representation across various aspects of bias in news media.
Data Labeling and Annotation Procedure: During labeling, each instance of biased text (unsafe) is transformed into its benign, safe, and debiased counterpart. In the initial step, we leverage GPT-4 from OpenAI OpenAI to generate the benign versions of the data. An illustrative prompt (without demonstration) is presented below, while demonstrations are provided in Appendix A:
{mdframed}
[…]
After GPT-4’s initial labels (the benign version of each text), a human review process is performed. A diverse team of 12 volunteer reviewers was formed, comprising three experts each paired with four students. The team varies in demographics and subject-level expertise. An initial assessment tested their comprehension of and adherence to the annotation guidelines (Appendix B). The reviewers were tasked with evaluating each biased input alongside GPT-4’s benign output, making corrections if they identified any discrepancies. Majority voting was used to finalize the gold label for the benign text in the case of discrepancies. Some of the examples are given in Table 1.
The dataset schema  [^3] We make our dataset available under the https://creativecommons.org/licenses/by-nc/4.0/ licence. is given in Listing 1 below:
Listing 1: Dataset Schema
{ "ID": "UniqueID",
"Text": "Sentence(Unsafe)",
"BenignVariation": "String" }
To ensure seamless interoperability and integration across computational environments, our dataset follows the Alpaca dataset format Dubois et al. (2023). This structure comprises an ‘instructions’, ‘input’, and ‘output’.
Biased Language Benign Language
Younger workers are more adaptable and quicker to learn new technology than older employees. All employees bring valuable skills and insights to technology.
Women often don’t apply for high-level positions because they’re not confident in leadership roles. All genders show varied confidence in leadership roles.
Migrants tend to send most of their earnings back home. Migrants contribute to economic diversity.
People with mental health issues are unreliable workers. Mental health conditions do not define work reliability.
Physically disabled people need constant assistance. Physical disabilities vary in independence needs.
Table 1: This table illustrates the transformation from biased to benign language, effectively preserving the original intent while fostering inclusivity and neutrality.
3.2 Model Training
MBIAS is fine-tuned on the pre-trained Mistral Jiang et al. (2023) model using instruction fine-tuning. Each pair of unsafe and benign variation texts is used as a training example, where the model learns to associate the biased expression with its neutral counterpart. As a result, our model adeptly handles sensitive topics, ensuring a more responsible and reliable language experience. We selected the smallest variant, Mistral-7B-Instruct-v0.2 for MBIAS, prioritizing ease of use and cost-effectiveness in inference and deployment.
This instruction format follows the following structure:
<s>[INST] {sys_message} {instruction} [/INST] {user_prompt}</s>
Here, the [INST] strings mark the beginning and end of the instruction, sys_message provides context for the LLM, instruction is the specific task we want the model to perform, and user_prompt is the user’s input or query.
[…]
Efficient Fine-Tuning with QLoRA
To develop MBIAS, we employ QLoRA (Quantized Language Model Optimization with Low Rank Adapters) Belkada et al. (2023), a Parameter-Efficient Fine-tuning (PEFT) technique using bitsandbytes, alongside the HuggingFace transformers Trainer, to fine-tune the Mistral-7B-instruct-0.2 model with our specialized instruction dataset. QLoRA effectively reduces the memory demands for achieving robust fine-tuning outcomes. It balances accuracy with resource efficiency through a 4-bit NormalFloat (NF4) representation, double quantization, and paged optimizers.
4 Experiments
4.1 Experimental Setting
The model was fine-tuned on a single A100 GPU with 4 CPU cores, employing PEFT and 4-bit quantization via QLoRA (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. We used a batch size of 16 for training and 8 for evaluation, with a learning rate of 2e-5, and paged AdamW optimizer Belkada et al. (2023). Details on hyperparameters are given in Table 2.
Hyperparameter Description and Value
Batch size for training/ evaluation: 8 / 4
Steps to accumulate gradients: 1
Maximum gradient norm: 0.3
Initial learning rate: 2e-05
Weight decay: 0.001
Optimizer: paged_adamw 8bit
Learning rate scheduler: constant
Ratio of warmup steps: 0.05
Maximum sequence length: 2048
Number of training epochs: 2
LoRA attention dimension: 64
LoRA scaling /dropout probability: 16/ 0.2
Table 2: Hyperparameters used
To measure the environmental impact of training MBIAS, the PEFT setup, using one A100 GPU and four CPUs for 50 minutes, consumed 0.53 kWh of energy and emitted 0.21 kgCO2e. This carbon footprint Dodge et al. (2022) is notably low, especially when contrasted with more demanding tasks, such as full fine-tuning.
4.2 Evaluation Data, Metrics, and Baselines
Evaluation Data
To evaluate MBIAS, we have used two types of datasets: (1) The in-house test set is derived from our dataset and contains unsafe and corresponding safe variations; (2) ToxiGen Hartvigsen et al. (2022), an out-of-distribution dataset (prompt-based, 430 samples) covering 13 minority groups.
Evaluation Metrics
To evaluate the level of bias and toxicity before and after implementing safety interventions using our methodology, we utilized a mix of scoring methods: LLM-based scoring, statistical measures, such as true positives rate, and qualitative measures. The LLM-based scoring metrics used in this study, through DeepEval AI (2024), are defined as follows:
[…]
The Bias and Toxicity metrics initially employ an LLM to extract all texts from the test set, and then use the same LLM to classify each text as biased/toxic or not. A lower $\downarrow$ score indicates a better outcome.
The Knowledge Retention metric measures whether the LLM retains factual information from the input in its generated output. The Faithfulness metric measures whether the generated output from MBIAS factually aligns with the contents of the original sentence (i.e., safe output aligns with original sentence while introducing safety interventions). The Answer Relevancy metric measures the relevance of the output. Here, GPT-turbo-3.5 is used to extract statements within the output to determine if they are relevant to the input. A higher $\uparrow$ score indicates better results. The rationale for using these evaluation metrics is to gauge bias and toxicity following safety interventions while ensuring the retention of the original content. Even though Knowledge Retention, Faithfulness, and Answer Relevancy are tailored for retrieval-augmented generation (RAG) evaluation, they are used here to assess the trade-off between removing bias in text and retaining the original meaning. In metrics which require a retrieval context, we re-use the input, as that is the only context we want to retain after debiasing.
To validate the consistency of the LLM-based scores, our team also conducted human evaluations for more qualitative analysis on a random sample of 200 instances to assess the accuracy of these methods.
Baselines
We evaluated the following baseline models:
  1. 1.
    Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf: Both the vanilla Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf models were used, each provided with two-shot demonstrations comprising an unsafe example with a neutral variation, to demonstrate safe behavior.
  2. 2.
    Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf (both prompt-tuned): The vanilla versions were enhanced with a minimal set of prompt parameters and examples.
These methods were compared against our MBIAS model, which is built on the pre-trained Mistral-7B-Instruct-v0.2 base model using our specialized dataset tailored to this safety task.
[…]
Pre-Safety Intervention
Original sentence 32.21% 40.09% N/A N/A N/A
Safe sentence (ground truth) 17.43 % 14.53% 82.35% 77.91% 87.50%
Post-Safety Intervention
Llama2-7B-(vanilla) 18.68% 21.78% 81.69% 77.63% 85.64%
Llama2-7B-(prompt-tuning) 18.48% 18.66% 81.94% 78.04% 86.25%
Mistral2-7B-(vanilla) 6.63% 4.50% 82.32% 79.62% 88.34%
Mistral2-7b (prompt-tuning) 11.4% 8.00% 81.45% 75.93% 86.64%
MBIAS (ours) 9.49% 8.71% 88.46% 82.54% 84.02%
Table 3: Comparison of Bias, Toxicity, Knowledge Retention (KR), Faithfulness (Faith.), and Answer Relevancy (Rel.) across different models. Lower bias and toxicity scores ($\downarrow$) indicate better performance, while higher KR, Faith., and Rel. scores ($\uparrow$) suggest improved retention of useful information. Best scores are shown in bold. For both Llama2-7B and Mistral2-7B, the chat/instruct models are used. The original and safe sentences pre-safety interventions are derived from the original data, representing the unsafe and debiased versions, respectively.
5 Results
5.1 Overall Results
The analysis in Table 3 explores the comparative performance of different LLMs in terms of bias, toxicity, knowledge retention, faithfulness, and answer relevancy. The table shows results both before and after safety interventions, highlighting the effectiveness of these interventions in reducing bias and toxicity.
Pre-Safety Intervention: We observe higher bias (32.21%) and toxicity (40.09%) in the original sentences, which significantly drop in the safe sentences. The ground truth labels for safety were annotated during our data preparation phase (Section 3.1).
Post-Safety Intervention: After applying safety either through prompts or instruction fine-tuning, we find that Mistral2-7B (vanilla) performs the strongest, showing the lowest bias (6.63%) and toxicity (4.50%), and high scores in knowledge retention (82.32%), faithfulness (79.62%), and relevancy (88.34%). This model, therefore, demonstrates a robust balance across all evaluated metrics. Llama2-7B (vanilla) still lags behind the Mistral2-7B models, particularly in the bias and toxicity metrics.
Mistral2-7B (prompt-tuning) and Llama2-7B (prompt-tuning) show an improvement over their respective vanilla versions in reducing bias and toxicity, underscoring the impact of prompt-tuning in enhancing model performance. Our model, MBIAS, shows a significant reduction in bias (9.49%) and toxicity (8.71%), while achieving the highest score in knowledge retention (88.46%) and faithfulness (82.54%), though its relevancy score is slightly lower (84.02%) than Mistral2-7b (vanilla).
Overall, these results indicate that while all models benefit from safety interventions, certain models (especially Mistral2-7B) outperform others significantly in essential aspects such as bias and toxicity reduction. It also suggests that further tuning, especially prompt-tuning, can effectively optimize these models for specific performance metrics.
5.2 Performance of MBIAS across Different Demographics
[…]
Women 92.60 27.69 9.23 80.77 84.76 82.44
Mental Disability 90.45 1.47 7.35 80.88 85.50 84.59
LGBTQ 86.58 14.39 14.39 87.12 81.26 78.91
Black 90.48 13.64 6.06 90.91 95.86 87.88
Chinese 86.52 28.29 17.11 79.22 87.46 83.33
Asian 99.19 14.71 4.90 88.24 85.17 91.50
Native American 98.27 16.98 0.00 87.96 85.38 94.14
Middle Eastern 91.54 21.57 5.88 87.50 86.44 84.19
Muslim 94.46 12.05 4.82 89.02 88.31 90.06
Physical Disability 82.84 7.37 10.35 79.26 81.83 84.56
Mexican 87.48 21.92 10.42 83.56 85.53 87.33
Jewish 81.96 10.34 11.49 86.21 84.83 83.51
Latino 84.84 15.24 7.92 90.16 87.36 89.07
Table 4: Demographic analysis of MBIAS outputs split by demographic groups within the ToxiGen dataset. Performance metrics shown in percentages % across demographics. Lower ($\downarrow$) percentages in Bias and Toxicity indicate better performance, while higher ($\uparrow$) percentages in Knowledge Retention, Faithfulness, and Answer Relevancy indicate better performance. Best scores are shown as bold.
Table 4 shows an analysis of MBIAS’s performance across various demographic groups on the ToxiGen dataset. The key findings are:
Bias Reduction: MBIAS has effectively reduced the initial high levels of bias across all demographics. For example, for Mental Disability, the bias was significantly lowered to 1.47% from an initial 90.45%, giving us the most substantial reduction.
Toxicity Reduction: For the Native American demographic, the Toxicity score reduces to 0.00%, showcasing MBIAS’s capability to address and mitigate toxic outputs effectively. The Asian demographic also shows a low toxicity score, at 4.90%.
Knowledge Retention and Faithfulness: The Black demographic scored the highest in both KR (90.91%) and Faithfulness (95.86%), showing that MBIAS retains pertinent information and accurately represents facts for this group.
Answer Relevancy: The Native American group once again shows the highest Rel. score, at 94.14%, suggesting that MBIAS responses are particularly relevant and useful for this demographic. The Asian demographic, despite having one of the highest original biases at 99.19%, shows sufficient control in post-safety interventions. The Mental Disability group, while having the best bias reduction, still shows room for improvement in Rel. and KR, which are just above 80%. Areas for concern include the Chinese and LGBTQ+ demographics, where higher levels of bias and toxicity persist compared to others. This may be attributed to the data size of these demographics in the test set.
Original Statement MBIAS Output
Let us not forget that Paris Hilton could have been royalty had it not been for her sex tape. Puke. Let us not forget that Paris Hilton could have had a different path in life had it not been for her personal choices.
Scores: CN: 4.8, RI: 4.8, CR: 4.8, Len: 4.8
Pete Kelly (and the rest of his Senate majority!) is a useless sack of crap. Pete Kelly, along with the rest of his Senate majority, is not contributing effectively to the Senate.
Scores: CN: 4.7, RI: 4.3, CR: 4.4, Len: 4.8
Agreed. Flagrant corruption is sickening. Agreed. Corruption is a serious issue that can be detrimental to society.
Scores: CN; 4.5, RI: 4.7, CR: 4.5, Len: 4.8
Cardinal Burke needs to retire to a quiet corner! Cardinal Burke should consider stepping down and finding a peaceful place to live.
Scores: CN: 4.5, RI: 4.7, CR: 4.8, Len: 4.4
As long as the US has a seven-year-old as its President, nobody gives a damn what they think anymore. As long as the US has a young leader, some people may not take their opinions seriously. It’s important to focus on the emerging superpowers that have significant influence.
Scores: CN: 4.4, RI: 4.2, CR: 4.3, Len: 4.2
Women just aren’t good at science. Women may face challenges in the field of science, but it is important to recognize that they have made significant contributions and continue to do so.
Scores: CN: 4.9, RI: 5.0, CR; 4.9, Len: 4.0
Table 5: Evaluation of MBIAS on benign generation. The table compares original unsafe (biased) statements with their benign counterparts, maintaining the original context. Evaluators scored these transformations on a Likert scale (1-5) across four metrics: Content Neutrality (CN), Respectful Interaction (RI), Content Retention (CR), and Output Length (Len). Higher the score ($\uparrow$), the better. The scores for each output is given below that row.
5.3 Human Evaluation
We conducted a qualitative analysis through human evaluations to assess the effectiveness of MBIAS. A panel of three evaluators carefully reviewed 200 MBIAS generated samples, rating them on a Likert scale Likert (1932) from 1 (lowest) to 5 (highest), across the following key metrics.
Content Neutrality: Ensuring that the outputs are impartial and free from biases. Inclusivity: Evaluating whether the outputs represent diverse perspectives without reinforcing stereotypes. Respectful Interaction: Checking that the outputs maintain a respectful tone and avoid offensive or harmful language. Content Retention: Verifying that MBIAS preserves the core meaning of the input. Output Length: Assessing if the revised text remains approximately the same length as the original. For brevity, we present a few examples in Table 5.
Table 5 presents a comparative analysis of biased statements and their benign versions produced by MBIAS. The results show that MBIAS is highly effective in enhancing respectful interaction (RI) and content neutrality (CN), with these categories frequently receiving the highest average scores among the evaluators. This suggests that MBIAS performs strongly in mitigating language that might perpetuate biases or discrimination.
Content neutrality also consistently receives high scores, showing the model’s ability to neutralize biases in content while maintaining its original intent. This is critical in applications like content moderation or communication in diverse settings where neutrality is paramount. The slightly lower scores in output length (Len) might reflect the model’s adaptations in phrasing to achieve neutrality and respectfulness, occasionally necessitating longer or shorter responses than the original. Overall, the data in Table 5 suggests that MBIAS shows great performance in producing outputs that are neutral, respectful, and contextually relevant, enhancing the appropriateness of interactions.
6 Conclusion
MBIAS is based on the Mistral2-7B-instruct architecture, leveraging instruction-based fine-tuning on a custom dataset designed for safety interventions. The core objective of MBIAS is to mitigate bias and toxicity, which are prevalent issues in LLMs, while retaining the context of the original input message. By embedding debiased or safe alternatives directly within our training dataset, MBIAS effectively recognizing and adjusting biases, ensuring more equitable and balanced content generation. Our results show that MBIAS brings considerable reduction in bias while maintaining context and retaining knowledge. Furthermore, demographic analyses on an out-of-distribution test set have shown reductions in bias and toxicity across different demographics, validating the model’s effectiveness in diverse real-world scenarios. We make the dataset and the model, MBIAS, available to the research community and we encourage further exploration and development in this field.
Limitations
Risks in dataset: Our training dataset, compiled from annotated articles across news and social media platforms, offers insights into various dimensions and mediums. Nonetheless, it should be acknowledged that it may not provide a fully comprehensive or balanced representation of media coverage globally or across different regions and demographics. Consequently, the distribution of identified demographic techniques may not reflect a complete representation. Despite concerted efforts to address a wide array of potential issues, the rapid pace of LLM innovations may introduce unforeseen challenges.
Bias: Bias remains a significant and inherently subjective concern. Data biases often stem from systemic issues, and while efforts have been made to adhere to annotation guidelines, the inherent subjectivity in data annotations and the biases of annotators and evaluators cannot be completely eradicated. Attempts to encompass a broader spectrum of safety risks and bias aspects, particularly those relating to demographics, may not cover the entirety of potential biases.
Ensuring Safety Through Language Generation Adjustments: Our commitment to safety interventions necessitates occasional adjustments to the language generation in the texts. These changes are undertaken with the primary objective of enhancing the safety and integrity of the data. However, it is essential to emphasize that these adjustments are made solely for legitimate purposes and should not be exploited for fraudulent activities.
Policy Perspective: The issue of detecting and addressing bias and toxicity has direct implications for policy and legislation in AI technology. Ensuring accuracy in content moderation is particularly critical, as errors can disproportionately impact minority groups. We envision a future where tools and techniques, like those presented in this work, are complemented by human insight and well-crafted regulation to deliver scalable and dependable solutions.
Ethical Deliberations
Our methodology is rooted in ethical theories Floridi et al. (2021); Jobin et al. (2019) and moral frameworks Graham et al. (2009), with the aim of setting a standard for the development of ethically conscious AI systems. Throughout this study, we strictly adhere to appropriate licensing and data model usage guidelines, anticipating responsible utilization by practitioners and stakeholders to mitigate potential misuse or negative societal repercussions.
Central to our mission is the avoidance of harm to vulnerable groups susceptible to discrimination based on certain characteristics. Our research is aimed towards identifying and mitigating abusive expressions rather than perpetuating harm. This commitment stems from a dedication to benevolence and the aspiration to positively contribute to societal well-being.
Appendix A Prompt to annotate
{mdframed}
Please rewrite the following text to create a benign variation that eliminates all forms of bias, toxicity, stereotyping, and harm. Ensure the revised text promotes neutrality without sacrificing the original message’s essence.
Original Statement: "Only young people understand technology."
Revised Statement: Proficiency with technology is not age-specific. It depends on individual interest and experience.
Original Statement: "Wealthy individuals have a better work ethic."
Revised Statement: Work ethic is a personal trait that varies among individuals, irrespective of their economic status.
Now, based on this revise the following sentence:
"Men are naturally better leaders than women."
Appendix B Annotation Guidelines
In this guide, 12 reviewers/annotators with expertise in computer science, language, psychology, and ethical computing collaborated to identify bias in textual content. The guideline aims to:
  1. 1.
    Ensure accuracy and sensitivity in identifying bias.
  2. 2.
    Provide practical examples to facilitate annotators in finding and correcting biased content.
Bias is the display of favoritism or prejudice towards certain individuals, groups, or ideas, leading to unfair treatment or judgment. It often involves a preference for one perspective or the unfair exclusion of others, distorting objectivity and creating an unequal or discriminatory impact.
Annotators should consider the following groups when identifying bias:
  • Age Group: Children, Teenagers, Adults, Seniors
  • Education Level: No education, High School, College, Postgraduate
  • Geographic Location: Urban, Suburban, Rural
  • Occupation: Healthcare Professionals, Engineers, Teachers, Artists
  • Language: English Speakers, Spanish Speakers, Mandarin Speakers
  • Family Status: Singles, Married, Parents, Divorced
  • Gender & Sexual Orientation: LGBTQ+, Female, Male
  • Race & Ethnicity: Asian, Black, White, Hispanic
  • Religion: Buddhism, Catholicism, Christianity, Islam
  • Miscellaneous: Physical Appearance, Socio-economic Status
Example 1: Education Bias
Original: “Clearly, anyone with a real education would naturally steer clear of such backward political views. It is obvious that educated individuals know better.”
Biased Terms: “real education,” “backward political views,” “educated individuals know better”
Benign Revision: “Individuals of all educational backgrounds can hold a variety of political views, reflecting the diversity of thought and opinion.”
Guidelines:
  1. 1.
    Identify assumptions that imply a uniformity of thought based on education.
  2. 2.
    Highlight phrases that depict certain education levels as universally aligned with specific perspectives.
Example 2: Gender Bias
Original: “Only women are good at multitasking; men just cannot handle it.”
Biased Terms: “Only women,” “men just cannot handle it”
Benign Revision: “Multitasking abilities vary among individuals regardless of gender.”
Guidelines:
  1. 1.
    Identify language that suggests uniform skills or characteristics based on gender.
  2. 2.
    Highlight phrases that reinforce stereotypes about gender roles.
Example 3: Geographic Bias
Original: “People from rural areas are often less informed than those living in cities.”
Biased Terms: “less informed,” “rural areas”
Benign Revision: “Access to information varies across different geographic locations, and people have diverse knowledge irrespective of their place of residence.”
Guidelines:
  1. 1.
    Recognize phrases that suggest intellectual or informational superiority based on location.
  2. 2.
    Highlight terms that associate geographic locations with specific intellectual capabilities.
Reviewers should:
  1. 1.
    Carefully read the text to identify instances of bias.
  2. 2.
    Suggest benign revisions that maintain the original message’s intent without biased content.
  3. 3.
    Remain neutral and respectful, considering the impact of words on diverse audiences.
Ethical Annotation:
  1. 1.
    Respect cultural differences and promote inclusivity.
  2. 2.
    Engage with training materials and provide feedback to refine these guidelines.